Bent Gregorian Technical Observing : 07 September 2009 UT

Observer: J. Hill (mtn)
Telescope Operator: D. Gonzalez Huerta
SW support: N. Cushing, M. Delapena, C. Biddick, J. Davis (Tucson)

Summary

Lost the first few hours to clouds. We spent the time power cycling all the computers in the control room.

Then we opened for one hour and got to test the x-flips (with and without). They work as expected.

Then we closed for clouds again.

Opened again for half an hour later. Verified that collimation model (LUCIFER) is OK at EL=25 deg. And fixed a couple bugs in John's transtar.cl script caused by AGW renumeration.

It was a remarkably productive night given that we were only open for 1.5 hours. No problems with the telescope at all while observing.

Preparations

Butch and Brian installed the spare pump for the chilled water/glycol.

Torsten and Jim sorted out the remaining GCS communication issues.

Details

Thunderstorms through the afternoon, so we are doing software tests and such.

02:15 pointto Vega

02:20 OSS is hanging. Called Paul. We are cycling power on M3 UMAC. (Instructions for remote power cycling on the Engineering wiki.)

02:47 OSS and M3 UMAC are happy now. But OSS is complaining about secondary limits. Paul is updating limits, and restarting OSS.

Computer Power Off Test

03:30 - 05:00 Software did the test of shutting down all the computers in the control room and restarting them. Torsten had RPC problems reconnecting to AGW, but in general in went very smoothly. They did find a few interesting issues. The agw-control computer is the acting router between GCS and AzCam (this was the problem last night). The IDL license server likes to be started manually.

Open

05:16 Open with medium to heavy cirrus clouds. Rain over Vail and Benson, but nothing showing upstream of us. T=8.7 degC, D=4.6 degC.

register IRTC

05:33 pointto M5_0850 ACQUIRE gs=0- failed
Resmsg[3] = couldn't move probe to position (SFP) 0.860029, -4.05317, focus: 35.2853 :AGw command status: AGW3: Requested position is Out of Range

05:37 pointto M5_0855 ACQUIRE gs=0 - succeeded by we didn't see a star.

05:39 pointto M5_0850 ACQUIRE gs=0 - succeeded but we didn't see a star initially. Found star at IE=-67 CA=-52 with ptspiral. (IRTC pointing model)

05:47 pointto BS9185 gs=0 ACQUIRE

05:48 pointto BS9185 gs=0 ACTIVE - guiding is OK

05:53 authorize LUCIFER

05:54 register LUCIFER

05:55 pointto M5_0850 ACQUIRE gs=0 - Found star at IE=-94 CA=-67 (Lucifer pointing model)

06:01 pointto BS9185 gs=0 ACTIVE - Torsten is pleased that the acquisition offset did not work, since he had not yet flipped the sign.

Torsten changes x-values in last 4 variables in AIP_L.cfg from 1 to -1
pointingCorr_flip_x     -1.0            float   # should the SFP X coordinate be flipped before sending to PCS for initial pointing correction? -1.0 means flip
pointingCorr_flip_y     1.0             float   # should the SFP Y coordinate be flipped before sending to PCS for initial pointing correction? -1.0 means flip
guidingCorr_flip_x      -1.0            float   # should the SFP X coordinate be flipped before sending to PCS for guide correction? -1.0 means flip
guidingCorr_flip_y      1.0             float   # should the SFP Y coordinate be flipped before sending to PCS for guide correction? -1.0 means flip

06:04 pointto BS9185 gs=0 ACTIVE - Now acquisition offset and guiding succeed on-axis.

Seeing is about 0.7 arcsec.

06:08 pointto StoneOagw TelM="ACTIVE" gs=1 - guiding and acquisition are OK.

06:12 pointto StoneOagw TelM="ACTIVE" gs=1 PA=330 (off the y-axis)

06:15 Clouds move over, so we lost the guide star. Somewhat of a test of guiding on faint stuff.

06:17 WFS is putting in big quantities of trefoil when the clouds make the guide star faint relative to the lunar background. (We've set to Doug to work thinking about how to supress this.)
JMH: This may have been caused by the x_pupil_shift problem that we discovered the following night.

Close

06:20 Close for clouds.

06:30 Michelle is switching us to a new version of PCS which has the "new" rotator handling for LBC.

09:00 The only significant cloud in Arizona seems to be sitting right above us.

Open

09:42 Re-open as the clouds are thinning out.

09:48 pointto WT10_221 TelM="ACQUIRE" PAM- gs=0 - star just at corner of the image. Remove the spurious secondary offload. Adjust IE=-111 CA=-94

09:55 pointto StoneOagw TelM="ACTIVE" gs=1 PA=330. This is at EL=26 deg. Uncorrected image is 1.8 arcsec. (Collimation model is still OK down here.) Image improves to 0.9 arcsec after a bit of focussing.

09:59 cl < tran_StoneO_00.cl

gs21 gs17 The transtar script is broken - complaining about invalid AGW unit number. So this is turning out to be a test of multiple presets rather than the IRTC data to check the transform that I was hoping for. Corrected the AGW unit check in transtar.cl, so we got a couple IRTC images at the end.

10:10 cl < tran_StoneO_00.cl again starting at EL=23 deg. failed to acquire gs21, and wanted a password for agwuser@lbtmu105 (the public key was only on agw-control)
JMH: I added a public key for LBTO@lbtmu42 to aqwuser@lbtmu105 the next afternoon.

GCS GUI reports that AGW unit is "DISCONNECTED" after moving the probe. What does that mean?

10:14 pointto StoneOagw ACQUIRE gs=0

but no star because it is clouds.

Close

10:16 Close for returning clouds. T=8.8degC, D=2.1 degC Wind light.

This marks the end of this night.

-- JohnHill - 07 Sep 2009
Topic revision: r6 - 08 Sep 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback